OpenStudy (anonymous):

is 2/5x=y a direct variation

OpenStudy (anonymous):

is this the equation: \(\large \frac{2}{5}x=y \) or is it this: \(\large \frac{2}{5x}=y \)

OpenStudy (lgbasallote):

\(\frac{2}{5x} = y\) is not a direct variation \(\frac{2}{5}x = y\) is a direct variation

OpenStudy (lgbasallote):

choose your poison ;)

OpenStudy (anonymous):

thank you all

OpenStudy (anonymous):

cyanide.

OpenStudy (lgbasallote):

potassium?

OpenStudy (lgbasallote):

or pure?